# When ControlNet Meets Inexplicit Masks: A Case Study of ControlNet on its Contour-following Ability

### v2

@@ 6.1 Implementation Details
We employ CLIP-Score and FID for evaluation. As CR metric would fail when the strict alignment of contours is violated, we propose two additional metrics inspired by  Hinz et al. (2020), namely Layout Consistency (LC) and Semantic Retrieval (SR). Definitions are as follows.
Definition 1 (Layout Consistency, LC).
@@
$LC=\frac{|({\bigcup_{i}^{M}a_{i}})\cap({\bigcup_{j}^{N}b_{j}})|}{|({\bigcup_{i}^{M}a_{i}})\cup({\bigcup_{j}^{N}b_{j}}|)}.$ (7)
Definition 2 (Semantic Retrieval, SR).
@@
$SR=\frac{|S\cap L|}{|L|}.$ (8)
Specifically, we utilize an open-vocabulary object detector, i.e., OWL-ViT Matthias Minderer (2022) following VISOR Gokhale et al. (2022). We set the confidence threshold $t=0.1$ and use instance labels as prompts.
@@
LC ControlNet-$\boldsymbol{m}_{0}$ 0.605 0.531 0.489 0.469 0.466 0.465 0.504
@@
SR ControlNet-$\boldsymbol{m}_{0}$ 0.522 0.440 0.374 0.327 0.320 0.303 0.381
Ours 0.513 0.503 0.495 0.482 0.477 0.462 0.489
Table 1: Performance comparison of our method and baselines, i.e., the vanilla ControlNet (ControlNet-$\boldsymbol{m}_{0}$) and ControlNet-$\boldsymbol{m}_{\infty}$ trained with bounding-box masks under different dilation radius $\boldsymbol{r}$.
6.2 Comparisons with Vanilla ControlNet
As Tab. 1 depicts, our method achieves competitive or better performance on all metrics, demonstrating robustness under all dilation radius. Visualized results in Fig. 7 further illustrate the effectiveness of our method. ControlNet-$\boldsymbol{m}_{0}$ tends to blindly follow the outlines of control inputs, thus suffering from unnatural images or violation of user intentions, especially when it is conditioned on largely deteriorated masks. On the other hand, ControlNet-$\boldsymbol{m}_{\infty}$ struggles with precise masks with fine details like misinterpreting a single instance mask into multiple objects. In contrast, our method excels in understanding object orientations and shapes according to inexplicit masks, thanks to additional guidance of shape priors. More examples are available in the Appendix.
6.3 Ablation Studies
@@
Table 2: Comparison with random dilation for augmentation.
@@
Table 3: Performance comparison between $\rho$ and predicted $\hat{\rho}$.
Comparison with random dilation augmentation.
As augmentation with random dilated masks serves as an intuitive and straightforward solution to improve the robustness of deteriorated masks, we compare this strategy with our Shape-aware ControlNet in Tab. 2. While this data augmentation strategy is also effective, our method achieves consistently higher performance on all metrics. Moreover, we highlight that our method provides users with the convenience of easily modifying the shape of generated objects through additional control over shape priors.
Accuracy of deterioration estimator.
We validate the accuracy of our proposed deterioration estimator in Fig. 8. The overall average $L1$ error is 5.47%. Owning to the robustness of the modulation block, such an error shows minimal impact on the fidelity of generated images as validated in Tab. 3. Detailed error curves are presented in the Appendix.
Figure 8: The error of deterioration estimator when radius $r=20$.
Robustness on shape priors.
We further examine the robustness of our method on different shape priors. We take dilation radius $r=20$ and manually adjust the shape prior $\rho$ to $(\rho+\Delta\rho)\in[0,1]$, where $\Delta\rho\in[0,1]$. As shown in Fig. 9, changing $\rho$ reveals little impact on the text faithfulness and image quality. But lower $\rho$ encourages better adherence to conditional masks, resulting in slightly higher LC with the assigned control masks, and vice versa. Visualized examples of tuning $\rho$ can be found in Fig. 11.
Figure 9: The variation of metrics under $(\rho+\Delta\rho)$ when $r=20$.
6.4 Applications
We showcase several applications of our method. For page limitations, additional examples are included in Appendix.
@@ 6.4 Applications
Since our Shape-aware ControlNet interprets inexplicit masks robustly, we observe its effectiveness in handling programmatic sketches and human scribbles as shown in Fig. 10. Our method generates reasonable images with abstract or inexplicit masks while keeping high fidelity and spatial control. We follow Zhang et al. (2023b) to prompt GPT-4 to produce TikZ codes for object sketches. Human scribbles are converted from Sketchy Sangkloy et al. (2016).
Figure 10: Performance comparison on TikZ sketches and scribbles.
@@
While the deterioration estimator provides the deterioration ratio $\hat{\rho}$ for reference, we can manually set $\rho$ to control the shape prior. It determines how much the generated objects conform to the provided conditional masks as shown in Fig. 11. We notice it is possible to extend $\hat{\rho}$ to values other values, even though $\rho\in[0,1]$ for training. A lower $|\rho|$ encourages stronger adherence to the contours of the provided masks.
Figure 11: Shape-prior control via deterioration ratio $\rho$.
@@
We realize composable shape control via a multi-ControlNet structure Zhang et al. (2023a), where we generate images by assigning different priors to each part of control masks. This enables strict shape control over specifically assigned masks while allowing T2I diffusion models to unleash creativity in imagining objects of diverse shapes within inexplicit masks. A vivid example is presented in Fig. 12, where the parrots share fixed contours but hats are generated within masked areas in various styles, and vice versa.
Figure 12: Examples of composable generation with explicit and inexplicit masks, showing flexible control with diverse results.
7 Conclusion
In this paper, we reveal insights into the core traits of ControlNet, i.e., the contour-following ability. We quantitatively validate this property by distorting conditional masks and tuning hyper-parameters, where we uncover severe performance degradation caused by inexplicit masks. In light of this, we propose a novel deterioration estimator and a shape-prior modulation block to endow ControlNet with shape-awareness to robustly interpret inexplicit masks with an extra shape-prior control. This exhibits the potential of employing ControlNet in more creative scenarios like user scribbles.
@@ 7 Conclusion
  • Additional visualized examples with detailed illustrations are presented to supplement the main paper. (§A)
  • We give a brief discussion about future works. (§B).
Appendix A Additional Qualitative Examples
@@
We provide additional qualitative results to supplement the main paper. Tab. 4 gives a quick overview of all figures presented in Appendix and indicates where they are referenced in the main paper for supplementary. Details are as follows.
Fig ID Reference (main paper) Brief Illustration
@@
Fig.13 §1, Line 100 More examples of ControlNet and T2I-adapter to illustrate their contour-following ability.
Fig.14 §4.2, Line 279 Visualization of the inductive bias of dilation radius $r$.
Fig.15 §4.3, Line 302, 339 Visual results for tuning hyperparameters.
Fig.16 §6.2, Line 443 More comparison examples of our method and ControlNet.
Fig.17 §6.3, Line 460 Error analysis for the deterioration estimator.
Fig.18 §6.4, Line 472 Generation with TikZ sketches.
Fig.19 §6.4, Line 472 Generation with user scribbles.
Fig.20 §6.4, Line 472 Examples for modifying shape priors.
Fig.21 §6.4, Line 472 Examples for composable shape-controllable generation.
Table 4: Quick overview of the figures in this Appendix.
Figure 13: More examples of ControlNet and T2I-adapter illustrating spatially controllable generation with masks of varying precision. Both methods exhibit strong preservation of contours during the generation process. Green denotes the recalled contours and blue denotes the missing ones. Moreover, inexplicit masks with inaccurate contours would cause severe degradation of image fidelity and realism.
Implication of dilation radius $r$.
@@
We further visualize examples of ControlNet-$\boldsymbol{m}_{r}$ with conditional masks of varying precision in Fig. 14. Note that ControlNet-$\boldsymbol{m}_{r}$ is trained with dilated masks $\boldsymbol{m}_{r}$. As depicted in Fig. 14, ControlNet-$\boldsymbol{m}_{r}$ implicitly assumes the presence of dilation in the provided masks, even for precise masks. For instance, the zebra generated by ControlNet-$\boldsymbol{m}_{r}$ with conditional mask $\boldsymbol{m}_{r}$ (in red boxes) shares similar size and shape with the precise mask $\boldsymbol{m}_{0}$, which explains the high CR scores in Fig.3 in our main paper. In contrast, our Shape-aware ControlNet robustly interprets all deteriorated conditions thanks to additional control over the shape priors.
Figure 14: Visualized examples to illustrate the inductive bias, where ControlNet-$\boldsymbol{m}_{r}$ implicitly learns the dilated radius $r$ from the data. Therefore, it always assumes a dilation in the provided mask. The ControlNet-$\boldsymbol{m}_{r}$ conditioned on $\boldsymbol{m}_{r}$ (in red boxes) produces objects with similar shapes and sizes of mask $\boldsymbol{m}_{0}$. While ControlNet-$\boldsymbol{m}_{\infty}$ also breaks the inductive bias, our method surpasses it in terms of image fidelity and additional control over the shape priors.
Visualized examples of tuning hyperparameters.
@@
Fig. 15 presents visualized results of the vanilla ControlNet under different hyperparameter settings, including the CFG scale $\omega$, conditioning scale $\lambda$, and condition injection strategy in our main paper. The CFG scale $\omega$ exhibits minimal impact on the contour-following effect but enhances image fidelity with higher saturation. Lowering the conditioning scale $\lambda$ and omitting conditions in the reverse sampling stages alleviate the strong contour instructions. However, because of the significant performance gaps between explicit and inexplicit conditions, bridging this gap and obtaining satisfactory results solely through hyperparameter tuning is challenging. Moreover, sudden appearance changes usually occur when we adjust hyperparameters, making it tricky to find the optimal setting. Therefore, it is necessary to develop a method to robustly interpret inexplicit masks.
Figure 15: Visualized examples under different hyperparameter settings, i.e., CFG scale, conditioning scale, and condition injection strategy. Though the CFG scale shows little impact on the contour-following ability, reducing conditioning scales and discarding conditions at early reverse sampling stages help to relieve contour instructions, resulting in better results with inexplicit conditional masks. However, it can be tricky and hard to achieve satisfactory results through hyperparameter tuning, especially for deteriorated control masks.
More visualized comparison with the vanilla ControlNet.
@@
In Fig. 16, we provide more examples of our Shape-aware ControlNet compared with the vanilla ControlNet. Our method not only achieves competitive results with ControlNet on precise masks, but also demonstrates an enhanced capability to interpret object shape and pose from inexplicit masks of varying precision.
Figure 16: More examples of our Shape-aware ControlNet compared with the vanilla ControlNet, i.e., ControlNet-$\boldsymbol{m}_{0}$ and ControlNet-$\boldsymbol{m}_{\infty}$.
Full results of error analysis for the deterioration estimator.
@@
Fig. 17 reports the full results of the estimation error for the proposed deterioration estimator. The overall averaged $L1$ error under different dilation radius $r$ is 5.47%. We confirm that this error has little impact on the performance due to the robustness of the shape-prior modulation block, as discussed in the main paper §6.3.
Figure 17: Error analysis of our deterioration estimator at different dilation radius $r$. The overall average $L1$ error is 5.47%. Notably, such errors show little impact on the performance of our Shape-aware ControlNet, referring to our main paper §6.3.
More application examples.
@@
We showcase additional examples of our method in three application scenarios, including robust generation with TikZ sketches (Fig. 18) following Zhang et al. [2023b], human scribbles (Fig. 19) converted from Sketchy Sangkloy et al. [2016], modification of shape priors (Fig. 20), and composable shape-controllable generation (Fig. 21). These examples verify that our method helps extend ControlNet to creative applications with more flexible control signals, owing to controllable shape priors.
Figure 18: More examples of our method on TikZ sketches compared with the vanilla ControlNet. Our Shape-aware ControlNet exhibits advanced performance on generation with abstract masks with inaccurate contours.
Figure 19: More examples of our method on human scribbles compared with the vanilla ControlNet. Our Shape-aware ControlNet exhibits advanced performance on generation with inexplicit masks with inaccurate contours.
Figure 20: More examples of controlling object shape via the deterioration ratio $\rho$. Despite training with $\rho\in[0,1]$, we empirically find it generalizes well to other values and offers additional control over the shape of generated objects. A small $|\rho|$ encourages generated objects to adhere more tightly to the outlines, and vice versa. Moreover, note that adjusting $\rho$ has little impact on image fidelity.
Figure 21: More examples of composable shape-controllable generation with our method. Leveraging a Multi-ControlNet structure, we can assign different priors to each part of the control masks. This enables strict shape control over specific masks, while simultaneously allowing T2I diffusion models to unleash creativity in imagining objects of diverse shapes within inexplicit masks.
Appendix B Future Works

### v3

@@ 6.1. Implementation Details
We employ CLIP-Score and FID for evaluation. As the proposed CR metric would fail when the strict alignment of contours is violated, we propose two additional metrics inspired by SOA (Hinz et al., 2020), namely Layout Consistency (LC) and Semantic Retrieval (SR). Definitions are as follows.
Definition 6.1 (Layout Consistency, LC).
@@
(7) $LC=\frac{|({\bigcup_{i}^{M}a_{i}})\cap({\bigcup_{j}^{N}b_{j}})|}{|({\bigcup_{i}^{M}a_{i}})\cup({\bigcup_{j}^{N}b_{j}}|)}.$
Definition 6.2 (Semantic Retrieval, SR).
@@
(8) $SR=\frac{|S\cap L|}{|L|}.$
Specifically, we utilize an open-vocabulary object detector, i.e., OWL-ViT (Minderer et al., 2022) following VISOR (Gokhale et al., 2022). We set the confidence threshold $t=0.1$ and use instance labels as prompts.
Table 1. Performance comparison with baselines, i.e., the vanilla ControlNet (ControlNet-$\boldsymbol{m}_{0}$) and ControlNet-$\boldsymbol{m}_{\infty}$ that trained with bounding-box masks, under different dilation radius $\boldsymbol{r}$. Our method exhibits advances in interpreting deteriorated masks and achieves the best average performance.
@@
LC ControlNet-$\boldsymbol{m}_{0}$ 0.522 0.440 0.374 0.327 0.320 0.303 0.381
Ours 0.513 0.503 0.495 0.482 0.477 0.462 0.489
SR ControlNet-$\boldsymbol{m}_{0}$ 0.605 0.531 0.489 0.469 0.466 0.465 0.504
@@
Table 2. Comparison with random dilation for augmentation.
6.2. Comparisons with the Vanilla ControlNet
As Tab. 1 depicts, our method achieves competitive or better performance on all metrics, demonstrating robustness under all dilation radius. Visualized results in Fig. 7 further illustrate the effectiveness of our method. ControlNet-$\boldsymbol{m}_{0}$ tends to blindly follow the outlines of control inputs, thus suffering from unnatural images or violation of user intentions, especially when it is conditioned on largely deteriorated masks. On the other hand, ControlNet-$\boldsymbol{m}_{\infty}$ struggles with precise masks with fine details like misinterpreting a single instance mask into multiple objects. In contrast, our method excels in understanding object orientations and shapes according to inexplicit masks, thanks to additional guidance of shape priors. More visualized examples are available in Appendix Fig. S9.
6.3. Ablation Studies
Comparison with random dilation augmentation.
As augmentation with random dilated masks serves as a straightforward solution to improve the robustness of deteriorated masks, we compare this strategy with our Shape-aware ControlNet in Tab. 2. Though data augmentation helps, our method outperforms it comprehensively on all metrics. Besides, our method supports shape control or variation by setting different ratios $\rho$ to control the alignment between conditional masks and synthesized images as presented in § 6.5. But the original ControlNet (even with random augmentation) fails.
Accuracy of deterioration estimator.
We validate the accuracy of the deterioration estimator. The average $L1$ error under all dilation radius is 5.47%. For page limitations, the error curves are presented in Appendix Fig. S10. Owing to the robustness of the modulation block, such a small error shows minimal impact on image fidelity as validated in Tab. 3. Discussions refer to Appendix §C.
Robustness on shape priors.
We further examine the robustness of our method on different shape priors. We take dilation radius $r=20$ and manually adjust shape prior $\rho$ to $(\rho+\Delta\rho)\in[0,1]$, where $\Delta\rho\in[0,1]$. As shown in Fig. 8, changing $\rho$ reveals little impact on the text faithfulness and image quality. But lower $\rho$ encourages better adherence to conditional masks, resulting in slightly higher LC with the assigned control masks, and vice versa. Visualized examples of tuning $\rho$ can be found in Fig. 10.
6.4. Extensive Results with More Conditions
It should also be noted that although the proposed method is designed for inexplicit masks, it can effortlessly adapt to other degraded conditions, demonstrating the versatility of our method in dealing with control signal deterioration. Here is an example with degraded edges and we leave other condition types for future works. For page limitations, quantitative results and visualizations are provided in Appendix §B. Specifically, we take deteriorated edges from masks in Eq. 4 for condition. The experimental settings are the same as § 6.1 and the only change is the condition used for training. Consistent with observations on object masks, the vanilla ControlNet blindly adheres to the provided deteriorated edges and struggles to generate high-quality images with inexplicit conditions. In contrast, our method adapts well to degraded edges of varying degrees, proving its effectiveness in dealing with other types of degraded conditions.
@@ 6.4. Extensive Results with More Conditions
Table 3. Performance comparison between $\rho$ and predicted $\hat{\rho}$.
Figure 8. The variation of metrics under $(\rho+\Delta\rho)$ when $r=20$.
Figure 9. Performance comparison between our method and ControlNet on TikZ sketches and scribbles.
Figure 10. Shape-prior control via deterioration ratio $\rho$.
Figure 11. Examples of composable generation with explicit and inexplicit masks, showing flexible control with diverse results.
6.5. Applications
Here we showcase several applications of our method. For page limitations, additional examples are included in Appendix §D.
@@ 6.5. Applications
While our model is only trained with dilated masks, we find it generalizes well to other types of realistic inexplicit masks and exhibits robust performance. Fig. 9 shows the effectiveness of our method in handling programmatic sketches and human scribbles. Our method generates reasonable images under abstract or inexplicit mask conditions while keeping high fidelity and spatial control. Here we follow Control-GPT (Zhang et al., 2023b) to prompt GPT-4 to produce TikZ codes for object sketches. Human scribbles are converted from Sketchy (Sangkloy et al., 2016) dataset.
@@
While the deterioration estimator provides the deterioration ratio $\hat{\rho}$ for reference, we can manually set $\rho$ to control the shape prior. It determines how much the generated objects conform to the provided conditional masks as shown in Fig. 10. We notice it is possible to extend $\hat{\rho}$ to a large range of other values, even though $\rho\in[0,1]$ for training. A lower $|\rho|$ empirically encourages stronger adherence to the contours of the provided masks.
@@
We realize composable shape control via a multi-ControlNet structure (Zhang et al., 2023a), where we generate images by assigning different priors to each part of control masks. This enables strict shape control over specifically assigned masks while allowing T2I diffusion models to unleash creativity in imagining objects of diverse shapes within inexplicit masks. A vivid example is presented in Fig. 11, where the parrots share fixed contours but hats are diverse in shapes, and vice versa.
7. Conclusion
This paper reveals several insights into the core traits of ControlNet, i.e., the contour-following ability. We quantitatively validate this property by distorting conditional masks and tuning hyper-parameters, where we uncover severe performance degradation caused by inexplicit masks. In light of this, we propose a novel deterioration estimator and a shape-prior modulation block to endow ControlNet with shape-awareness to robustly interpret inexplicit masks with an extra shape-prior control. This exhibits the potential of employing ControlNet in more creative scenarios like user scribbles and diverse degraded condition types.
Acknowledgements.
This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFC2705700, in part by the National Natural Science Foundation of China under Grants U23B2048, 62076186 and 62225113, and in part by the Innovative Research Group Project of Hubei Province under Grant 2024AFA017. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
@@ Acknowledgements.
  • Implementation details about the model structure and its computational complexity. (§ A)
  • Additional results on other kinds of deteriorated conditions, i.e., quantitative results on degraded edges. (§ B)
  • More discussions are provided, including the effects of imprecise $\hat{\rho}$, how to control the contour-following ability of our model, the advantages of inexplicit masks, and the difference between our work and Layout-to-Image generation. (§ C)
@@
  • Additional visualized examples with detailed illustrations are presented to supplement the main paper. (§ D)
  • We give a brief discussion about future works. (§ E)
Figure 12. The detailed structure of our proposed (a) deterioration estimator and (b) shape-prior modulation block.
Appendix A Details on Model Structure and Computational Complexity
Structures of the proposed modules. To illustrate our model design, we provide details on the structure of our deterioration estimator and shape-prior modulation block. As shown in Fig.12, the deterioration estimator is constructed by stacking 8 Conv+BN+SiLU units with 2 Linear+BN units. The shape-prior modulation block contains 12 two-layer MLPs with SiLU for each UNet’s encoder feature. Visit our github for more implementation details.
Model size and computational complexity. Tab. 4 reveals the model size computational complexity of the deterioration estimator and shape-prior modulation block. Our method introduces only marginal increments on parameters and computation complexity to realize shape-controllable generation, where it brings 35.09M (9.7%) additional parameters and 37.89G (16.2%) FLOPs in total compared with the raw ControlNet (361.28M Params, 233.17G FLOPs).
Table 4. Model size and computational complexity of our proposed Shape-aware ControlNet.
Method Params/M FLOPs/G
ControlNet 361.28 233.17
+shape-prior modulation +12.19 +0.02
+deterioration predictor +22.90 +37.87
Ours 396.37 271.06
Appendix B Quantitative Results with Deteriorated Edges
Though we mainly focus on one representative condition, i.e., object masks, in the main paper for illustration, our method can be applied to other kinds of degraded conditions with little effort as discussed in the main paper § 6.4. As shown in Fig.13, the vanilla ControlNet keeps adhering to the provided deteriorated edges blindly and struggles to generate high-quality images with inexplicit conditions, which are consistent with the observations on conditional masks. In contrast, our proposed method adapts well to degraded edges of varying degrees, proving its effectiveness in dealing with other types of degraded conditions.
We also provide quantitative results to further prove the effectiveness of our method on more condition types, i.e., deteriorated edges. Tab. 5 reports the CLIP-score, FID, Layout Consistency (LC), and Semantic Retrieval (SR) following the same setting as the main paper § 6.1. Compared with the vanilla ControlNet, our method not only obtains competitive results with accurate edges but also exhibits robust performance with degraded edges of varying degrees. These results demonstrate the advance of our method in handling diverse kinds of inexplicit conditions. More types of conditions are left for future works.
Table 5. Performance comparison of our method and the vanilla ControlNet, under the condition of degraded edges of different dilation radius $\boldsymbol{r}$. The results prove the effectiveness of our method in other kinds of degraded conditions besides masks.
Metric Method $\boldsymbol{r}$ AVG
0 20 40 80 100 $\infty$
CLIP- Score ControlNet 26.77 26.22 25.77 25.53 25.49 25.49 25.88
Ours 26.83 26.85 26.88 26.85 26.87 26.87 26.86
FID ControlNet 13.51 16.54 19.45 21.84 22.50 22.97 19.48
Ours 13.80 14.52 15.05 15.84 15.92 16.01 15.19
LC ControlNet 0.521 0.429 0.361 0.310 0.299 0.279 0.367
Ours 0.510 0.494 0.476 0.456 0.452 0.434 0.509
SR ControlNet 0.607 0.533 0.496 0.477 0.474 0.469 0.470
Ours 0.599 0.582 0.572 0.563 0.563 0.556 0.572
Figure 13. Comparison of ControlNet and our method with deteriorated edges. Our method can effortlessly adapt to diverse kinds of deteriorated conditions.
Appendix C More Discussions
The effects of imprecise estimation of $\rho$. As presented in the main paper § 6.3, the predicted deterioration ratio $\rho$ has an average error of 5.47%. While some readers may consider this error a little large, we argue that it is still within the tolerance of our method and does not mean the estimator is useless. As Fig. 14 shows, the model produces plausible and natural polar bears when $\rho=0.9$ (the ground truth $\rho\approx 1.0$), showing certain robustness for $\rho$. But a larger error over 50%, e.g. when $\rho=0.1,0.5$, still influences the image fidelity (the contour of the polar bear’s hip becomes wired), especially for relatively rough masks. Thus, considering the model’s robustness, a 5.47% average error of the predicted $\rho$ is acceptable, which means our estimator is still useful to provide a relatively accurate $\rho$ for reference. Besides, the deterioration ratio $\rho$ is a crucial value in our method, which controls a trade-off between spatial alignment with conditions and image fidelity. While the ground truth $\rho$ is accessible for training, $\rho$ is usually unknown in real-world applications during the inference period and varied for images. Thus, our deterioration estimator provides a solution estimating a suitable $\rho$ automatically. We will work on improving its accuracy in future works.
When does the model follow the given contours and when not? It is the shape-prior $\rho$ that decides whether the model follows the contour precisely or not. As shown in Fig. 14, when $\rho=0$, it means the conditional mask is precise, so the generated images will strictly follow the given masks like the original ControlNet. As $\rho$ grows larger, the model would pay more attention to the object position rather than exact contours. We provide two options to provide a reasonable $\rho$: 1) let the deterioration estimator automatically predict $\rho=\hat{\rho}$ based on the conditional inputs and prior knowledge, and 2) if needed, it allows to manually set a pre-defined $\rho=\rho_{0}$ to control how much the synthesized image follows the condition, e.g. set $\rho_{0}=0$ to follow the given contours.
Figure 14. The synthesized results with different $\rho$. The second row indicates the contour alignment between synthesized images and conditional masks.
Figure 15. Examples of generation with bounding boxes and inexplicit masks with different degrees of details.
Advantages of conditional inexplicit masks for shape-controllable generation. Considering that the bounding-box masks are easy to access, readers may be curious about the advantages of inexplicit masks, especially when masks are relatively rough. In other words, when it is better to use inexplicit masks than bounding-box masks? Here are some further explanations.
If the mask is relatively rough (with few details) and only prompts the object’s position, there will be no obvious distinctions in generation compared to that using bounding-box masks. In this case, we recommend using bounding-box masks for convenience. However, for some masks like sketches or scribbles, despite their roughness, they can still provide valuable information on details like facing direction, which helps to achieve more controllable results than a simple bounding box for creation. As shown in Fig. 15, take ‘drawing a horse‘ for example. If you only want to generate the horse at a certain position without other concerns, assigning a bounding box is the simplest way. However, when you want to control some details further like its facing direction or pose, using a rough mask containing some lines to indicate the head and legs will provide more satisfactory results than a simple bounding box.
Figure 16. Comparison with L2I method, i.e., LayoutDiffuse (Cheng et al., 2023). While our method achieves spatial control over the global layout, it may confuse the fine-grained layout for each object, which differs from the Layout-to-Image generation tasks.
Difference with the Layout-to-Image Generation. One may notice that when the dilation radius $r=\infty$, our Shape-aware ControlNet addresses a similar problem to Layout-to-Image (L2I) generation. But we claim that there exist subtle differences between L2I and our work. For a better understanding, explanations are as follows.
Layout-to-Image generation works, such as GLIGEN (Li et al., 2023), LayoutDifuse (Cheng et al., 2023), LayoutDiffusion (Zheng et al., 2023), etc., aim to generate objects conforming to the pre-defined layouts. Generally, L2I assigns fine-grained object positions in a tuple format of $(x,y,width,height,label)$, thereby associating each object to the specific bounding box. However, in our setting, only prompts and binary masks are available, which makes it hard to achieve fine-grained position control over each object like L2I. Fig. 16 compares our method with LayoutDiffuse (Cheng et al., 2023) for illustrations. While our method adheres to the global layout provided by bounding-box masks, the exact positions of each object are not fixed. A recent work (Lukovnikov and Fischer, 2024) explores adapting ControlNet to L2I generation, thus it is promising to extend our method to fine-grained layout control. However, as this is not the main topic of this paper, we leave it for future work. The main purpose of this paper is to examine the contour-following ability of ControlNet, and we reveal a solution to adapt ControlNet to deteriorated masks at varying precision levels.
Appendix D Additional Qualitative Examples
@@
We provide additional qualitative results to supplement the main paper. Tab. 6 gives a quick overview of all figures presented in the Appendix and indicates where they are referenced in the main paper for supplementary. Details are as follows.
ID Reference (main paper) Brief Illustration
@@
Fig.17 §1 More examples of ControlNet and T2I-adapter to illustrate their contour-following ability.
Fig.18 §4.2 Visualization of the inductive bias of dilation radius $r$.
Fig.19 §4.3 Visual results for tuning hyperparameters.
Fig.20 §6.2 More comparison examples of our method and ControlNet.
Fig.21 §6.3 Error analysis for the deterioration estimator.
Fig.22 §6.5 Generation with TikZ sketches.
Fig.23 §6.5 Generation with user scribbles.
Fig.24 §6.5 Examples for modifying shape priors.
Fig.25 §6.5 Examples for composable shape-controllable generation.
Table 6. Quick overview of additional examples.
@@
Implication of dilation radius $r$. We further visualize examples of ControlNet-$\boldsymbol{m}_{r}$ with conditional masks $\boldsymbol{m}_{r}$ of varying deterioration degrees in Fig. 18. Note that ControlNet-$\boldsymbol{m}_{r}$ is trained with dilated masks $\boldsymbol{m}_{r}$. As depicted in Fig. 18, ControlNet-$\boldsymbol{m}_{r}$ implicitly assumes the presence of dilation in the provided masks, even for precise masks. For instance, the zebra generated by ControlNet-$\boldsymbol{m}_{r}$ with conditional mask $\boldsymbol{m}_{r}$ (in red boxes) shares similar size and shape with the precise mask $\boldsymbol{m}_{0}$, which explains the high CR scores in Fig.3 in our main paper. In contrast, our Shape-aware ControlNet robustly interprets all deteriorated conditions thanks to additional control over the shape priors.
@@
Visualized examples of tuning hyperparameters. Fig. 19 presents visualized results of the vanilla ControlNet under different hyperparameter settings, including the CFG scale $\omega$, conditioning scale $\lambda$, and condition injection strategy in our main paper. The CFG scale $\omega$ exhibits minimal impact on the contour-following effect but enhances image fidelity with higher saturation. Lowering the conditioning scale $\lambda$ and omitting conditions in the early reverse sampling stages alleviate the strong contour instructions. However, because of the significant performance gaps between explicit and inexplicit conditions, bridging this gap and obtaining satisfactory results solely through hyperparameter tuning is challenging. Moreover, sudden appearance changes usually occur when we adjust hyperparameters, making it tricky to find the optimal setting. Therefore, it is necessary to develop a method to robustly interpret inexplicit masks.
@@
More visualized comparison with the vanilla ControlNet. In Fig. 20, we provide more examples of our Shape-aware ControlNet compared with the vanilla ControlNet. Our method not only achieves competitive results with ControlNet on precise masks but also demonstrates an enhanced capability to interpret object shape and pose from inexplicit masks of varying deterioration degrees.
@@
Full results of error analysis for the deterioration estimator. Fig. 21 reports the full results of the estimation error for the proposed deterioration estimator. The overall averaged $L1$ error under different dilation radius $r$ is 5.47%. We confirm that this error has little impact on the performance including CLIP-score and FID due to the robustness of the shape-prior modulation block, as discussed in the main paper § 6.3.
@@
More application examples. We showcase additional examples of our method in three application scenarios, including robust generation with TikZ sketches (Fig. 22) following Control-GPT (Zhang et al., 2023b), human scribbles (Fig. 23) converted from Sketchy (Sangkloy et al., 2016) dataset, modification of shape priors (Fig. 24), and composable shape-controllable generation (Fig. 25). These examples verify that our method helps extend ControlNet to creative applications with more flexible control signals, owing to controllable shape priors.
Appendix E Future Works